Roland Stigge Debian Developer address placeholder stigge@debian.org

Roland Stigge, Debian Project

University of Leeds Computer Based Learning Unit LEEDS, LS2 9JT United Kingdom

3. Januar 2004

Subject: The license of L^{A}TEX2HTML

Dear officials at Leeds University,

I’m a developer of the Free Software / Open Source project called Debian^{1} and personally maintain the quite popular package L^{A}TEX2HTML^{2} for it. This conversion tool, designed to translate from the L^{A}TEX publishing format to HTML, was initially written in 1993 by Nikos Drakos during his work at Leeds University. Its license is:

LaTeX2HTML License
==================
Original Copyright notice:
LaTeX2HTML by Nikos Drakos <nikos@cbl.leeds.ac.uk>
Some portions of this package are published under the
GNU public license. These are clearly marked in the header.
****************************************************************
General License Agreement and Lack of Warranty *****************
****************************************************************
This software is distributed in the hope that it will be useful
but WITHOUT ANY WARRANTY. The author(s) do not accept responsibility
to anyone for the consequences of using it or for whether it serves
any particular purpose or works at all. No warranty is made about
the software or its performance.
Use and copying of this software and the preparation of derivative

1Debian is one of the most famous Debian GNU/Linux Computer Operating System Distributions, see http://www.debian.org/ 2http://www.latex2html.org/

works based on this software are permitted, so long as the following
conditions are met:
o  The copyright notice and this entire notice are included intact
   and prominently carried on all copies and supporting documentation.
o  No fees or compensation are charged for use, copies, or
   access to this software. You may charge a nominal
   distribution fee for the physical act of transferring a
   copy, but you may not charge for the program itself.
o  If you modify this software, you must cause the modified
   file(s) to carry prominent notices (a Change Log)
   describing the changes, who made the changes, and the date
   of those changes.
o  Any work distributed or published that in whole or in part
   contains or is a derivative of this software or any part
   thereof is subject to the terms of this agreement. The
   aggregation of another unrelated program with this software
   or its derivative on a volume of storage or distribution
   medium does not bring the other program under the scope
   of these terms.
This software is made available AS IS, and is distributed without
warranty of any kind, either expressed or implied.
In no event will the author(s) or their institutions be liable to you
for damages, including lost profits, lost monies, or other special,
incidental or consequential damages arising out of or in connection
with the use or inability to use (including but not limited to loss of
data or data being rendered inaccurate or losses sustained by third
parties or a failure of the program to operate as documented) the
program, even if you have been advised of the possibility of such
damages, or for any claim by any other party, whether in an action of
contract, negligence, or other tortious action.

While this license was unquestionably written as a Free Software / Open Source license, the Debian project considers it “non-free” because of the second clause:

o  No fees or compensation are charged for use, copies, or
   access to this software. You may charge a nominal
   distribution fee for the physical act of transferring a
   copy, but you may not charge for the program itself.

This violates our self-imposed “Debian Free Software Guidelines” (DFSG^{3}), because it doesn’t allow the software to be sold. Please note that this unveils a quite ironical condition because Debian is one of the least commercial system distributions (not even directly providing physical CDs, DVDs or support contracts). Instead, it’s a wholly voluntarily driven project as e.g. Linux itself. But please also note that it lives from its users who have to aquire the software somehow. This is usually done by CDs and DVDs which are in turn sold by different vendors.

3http://www.de.debian.org/social contract#guidelines

I would like to inform you that the Debian Project has concluded the above quoted license as “non-free”^{4} and therefore removed it from it’s main distribution. To facilitate the already existing large user base of the program (and to move the package back into our main distribution), I would like to ask you for the permission to distribute the software under the protection of the GNU General Public License (GPL)^{5}.

In the following I will list some further reasons for doing so:

• The original author of the package, Nikos Drakos, has signalled to possibly agree to a license change in the described direction. But he wondered if he needs support by you, his former employer where he worked during the birth of L^{A}TEX2HTML. Ross Moore, the current maintainer of L^{A}TEX2HTML, also showed the willingness to cooperate.

• The GPL is a quite popular license in the area of Free Software / Open Source. In contrast to individual licenses like the current one for L^{A}TEX2HTML, it’s well known in the whole world and under constant review by the user and developer base to prevent legal problems (possibly like the one in question).

• As noted in the above quoted license itself, some parts of L^{A}TEX2HTML are already issued under the protection of the GPL. This probably means that the whole package should be licensed under the GPL anyway.

• Debian is not the only project distributing L^{A}TEX2HTML. Other distributions of the popular GNU/Linux system like RedHat and SuSE and different BSD derivates also use L^{A}TEX2HTML and probably have the same problem.

• The change would ensure the investment of Leeds University into the Free Software community by saving its further widespread use.

I’m looking forward to get an answer from you. Thanks in advance.

Sincerely,

Roland Stigge Debian Developer

4http://lists.debian.org/debian-legal/2003/debian-legal-200310/msg00383.html and follow-ups 5http://www.gnu.org/licenses/gpl.html